40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 10403     Entry time: Fri Aug 15 17:24:44 2014
Author: Harry 
Type: Update 
Category: General 
Subject: Fiber Temp. 

 Earlier today Q and I somewhat resurrected my old PER measurement setup so I could run the temperature characterization experiment.

Unfortunately, when I tried to use the fiber illuminator, no light came from the other end, causing me to fail my primary goal for the summer of "don't break anything." The fiber has been re-spooled and labeled appropriately. Also sorry.

In addition to this, Q and I scavenged parts from the telescopes on the PSL and Y End tables, which were either not functional, or needed to have their mode matching adjusted, since we're using the non-PM fibers for FOL, which have a different numerical aperture, and thus slightly different output modes.

Specifically, this is involved removing the rotational mounts, and appropriate beam dumping.

My "calorimeter" still remains intact, in case anyone wants to make this measurement in the future, as this is my last day in the lab.

It's also effective at keeping drinks cold, if you'd rather use it for that.

ELOG V3.1.3-